### v2

@@ 3.2. DSL Analysis
For the analysis of the identified security DSLs, we followed a four-step procedure.
1. Definition of aspects: For each research question, we defined relevant aspects according to which we analyzed the identified security DSLs. We outlined these aspects already when discussing the research questions in Section 1 and provide detailed reasoning for each of them in Section 4.
2. Free text responses: To describe each DSL according to these aspects, we first searched the corresponding papers from our dataset for relevant information to describe the DSL according to each aspect. If we did not find the information in the paper, or if the aspect was open-ended, e.g., whether there are practical applications of the DSL, we searched for further publications on the DSL and for websites or documentation of the DSL.
In this first step, we collected short free-text responses to each aspect. To do this, we divided the selected papers between the first two authors of this paper. While collecting the information on the security DSLs, the two authors were continuously synchronizing to discuss interpretations of single aspects for specific DSLs and to avoid divergence in the collected information.
3. Categorization: After collecting information for each security DSL, we discussed all captured free texts and derived categories from them. Since not all captured information was easily mappable to the identified categories and to improve data quality, the categorization was followed by a data sanitation step in which the first two authors this time divided the individual aspects between them and went through all captures free text answers as well as the full papers again to assign every DSL, depending on the aspect, one or multiple of the identified categories. Again, we supported this procedure with continuous discussions between the first and second author.
4. Statistical analysis: Finally, after classifying all DSLs according to the categories identified for each aspect, we statistically analyzed the captured data. To do this, we created statistics on the distribution of each category for each aspect but also calculated correlations among the different aspects. We present these outcomes in detail in Section 4 and discuss implications in Section 5.
@@ 4. Systematic Literature Review
In our SLR, we identified 120 security DSLs which we analyzed in depth to answer the RQs presented in Section 1. In what follows, we present the data extracted by this analysis.
@@ 4.1. RQ1: What Security DSLs have been presented in the scientific literature?
ASL 1997 generic modeling of access control policies AC D, C (Jajodia et al., 1997)
Bro (Zeek) 1999 detection of network attackers by passive monitoring ID C, RT (Paxson, 1999) (Sharma et al., 2024)
AdeLe 2001 model a database of known attack scenarios TM, ID C, RT (Michel and Mé, 2001)
Ponder 2001 generic modeling of access control policies AC D, C (Damianou et al., 2001)
SPL 2001 single description of global security policies AC D, C (Ribeiro et al., 2001)
@@
CORAS 2002 risk assessment based on semiformal system model TM R, D (Lund et al., 2011; Fredriksen et al., 2002) (SINTEF, 2014)
@@
SecureUML 2002 modeling access control policies and their integration AC D, C, RT (Basin et al., 2011; Lodderstedt et al., 2002; Basin et al., 2006, 2007) (Doser, 2007)
STATL 2002 intrusion detection of known attack scenarios ID C, RT (Eckmann et al., 2002)
UMLsec 2002 planning of secure system design R D (Jürjens, 2002, 2001) (Jürjens et al., 2024)
FlowCAML 2003 secure information flow language IFC C (Pottier and Simonet, 2003) (Simonet, 2003)
uniframe 2003 security contracts for component descriptions AC D (Burt et al., 2003)
Ray et al. 2003 composition of access control frameworks AC D (Ray et al., 2003)
Rei 2003 policies applicable to pervasive environments AC C, RT (Kagal et al., 2003; Kagal, 2002)
XACML 2003 common language for expressing security policies AC C, RT (Ferdous and Poet, 2014; Lorch et al., 2003; Parducci and Lockhart, 2017) (OASIS, 2021)
ESRML 2004 enterprise information security requirements for IS0-17799 R, AC R (Roy et al., 2004)
SAML 2005 exchanging authentication and authorization data AC C, RT (Ferdous and Poet, 2014; Hughes et al., 2005) (OASIS, 2005)
X-GTRBAC 2005 expressive specification of RBAC AC D (Bhatti et al., 2006, 2005a, 2005b, 2008)
WS-CoL 2005 policy assertions of user requirements for web services R C, RT (Baresi et al., 2005)
UMLintr 2006 specification of intrusion detection ID R, D, C, RT (Hussein and Zulkernine, 2006)
WS-SecurityPolicy 2006 security constraints and requirements for web services R C, RT (Bajaj et al., 2006; Nadalin et al., 2009) (OASIS, 2012)
AsmLSec 2007 specifying attack scenarios ID T (Raihan and Zulkernine, 2007)
OSL 2007 specify usage control policies AC D, C (Hilty et al., 2007)
SECTET 2007 policies for secure interorganizational workflows R, AC R, D (Haley et al., 2006; Alam et al., 2006, 2007b, 2007a; Breu et al., 2007)
SecureTROPOS 2007 identification and planning of security requirements R R, D (Mouratidis et al., 2005; Mouratidis and Giorgini, 2007; Matulevicius et al., 2012) (Mouratidis, 2015)
SMCL 2007 secure multi party computation CR C (Nielsen and Schwartzbach, 2007)
@@
Rodríguez et al. 2007 embed security requirements into business processes R R, D (Rodríguez et al., 2007)
WS-Policy4MASC 2007 specification of monitoring and control policies R D, C (Tosic et al., 2007)
CA-RBAC 2008 context-based access control AC C, RT (Kulkarni and Tripathi, 2008)
SELinks/FABLE 2008 enforcement of security policies in web applications AC, IFC C (Afshar and Fallah, 2011; Swamy et al., 2008) (Corcoran et al., 2008)
SFDL 2.0 2008 secure multi party computation CR C, RT (Ben-David et al., 2008; Malkhi et al., 2004) (Tang et al., 2015)
Breu et al. 2008 guide security investment decisions TM R, D (Breu et al., 2008)
USE 2008 specify and verify security policies AC D, C, RT (Sohr et al., 2008)
Cryptol 2009 functional correctness of cryptography CR C (Erkök and Matthews, 2009) (Galois Inc., 2024)
Fabric 2009 secure information flow in distributed systems IFC C (Liu et al., 2009, 2017) (Myers et al., 2019)
Nego-UCONABC 2009 access control for cloud services AC D (Chen et al., 2009)
ReAlSec 2009 identify countermeasures for known attack patterns TM T, DP (Hamdi et al., 2009)
S-Promela 2009 generic security policies AC C (Abbassi and El Fatmi, 2009)
Security Policy Model 2009 generating security policies from security requirements R, AC R, D, C, RT (Wolter et al., 2008; Menzel et al., 2009; Wolter et al., 2009)
SPDL 2009 simplifying the specification of SELinux security policies AC C, RT (Nakamura et al., 2009) (Ynakam, 2008)
Pavlich-Mariscal et al. 2010 access control policies AC D, C (Pavlich-Mariscal et al., 2010)
ADTrees 2010 specify and analyze attack-defense scenarios TM D (Kordy et al., 2010)
ASLan++ 2010 security goals of distributed system designs AC D (von Oheimb and Mödersheim, 2010)
CapDL 2010 capability-based access control AC D, RT (Kuz et al., 2010) (McLeod et al., 2024)
CRePE 2010 fine-grained security policies for android applications AC C, RT (Conti et al., 2010)
SecPAL 2010 simple but expressive, declarative authorization language AC C (Becker et al., 2010)
SSMF 2010 identification of possible attack vectors TM R (Li et al., 2010)
TASTYL 2010 secure multi party computation CR C, RT (Henecka et al., 2010)
X-Policy 2010 dynamic execution permissions for collaborative systems AC D (Qunoo and Ryan, 2010)
Caisson 2011 secure information flow language IFC C (Li et al., 2011)
cPLC 2011 generate secure code for cryptographic two-party protocols CR D, C (Bangerter et al., 2011)
Asnar et al. 2011 identification of permission-related threats R, TM R (Asnar et al., 2011)
UACML 2011 access control modeling language independent of model AC D (Slimani et al., 2011)
Bain et al. 2011 high-level code for secure execution platforms (MPC & FHE) CR C (Bain et al., 2011)
VALID 2011 security goals for virtualized environments R DP, RT (Bleikertz and Groß, 2011)
PTaCL 2012 security policies based on attribute-based access control AC D (Crampton and Morisset, 2012)
Hoisl et al. 2012 secure object flow in business process IFC D (Hoisl et al., 2014)
STRoBAC 2012 spatial temporal role-based access control AC D, C, RT (Thi et al., 2012)
@@
TTCN-3 2012 generation of penetration tests with attack patterns TM T (Stepien et al., 2012) (ETSI, 2023)
Mitchell et al. 2012 secure execution platform based on MPC or FHE CR C (Mitchell et al., 2012)
BCDSDL 2013 network availability with respect to botnet attacks IDE D (Huan et al., 2013)
context-aware RBAC 2013 context-aware access control policies AC D (Schefer-Wenzl and Strembeck, 2013)
CIL 2013 specification of access control policies for Android AC C, RT (Bugiel et al., 2013) (SELinux Project, 2023)
@@
CySeMoL 2013 probabilistic vulnerability assessment TM D, C (Holm et al., 2015; Sommestad et al., 2013) (KTH Network and Systems Engineering Group, 2020)
DJS 2013 solution for untrusted code in javascript CR C (Bhargavan et al., 2013)
FAL 2013 secure logging configurations CR C, RT (Zawoad et al., 2013)
Paragon 2013 practical information flow control IFC C (Broberg et al., 2013a) (Broberg et al., 2013b)
Table 2. Security DSLs identified in the SLR (Part 2/2). Categories: Access Control (AC), Cryptography (CR), Information Flow Control (IFC), Intrusion Detection (ID), Requirements (R), Threat Modeling (TM); SDLC Phases: Requirements (R), Design (D), Code (C), Test (T), Deployment (DP), Runtime (RT)
@@
RBAC DSL 2013 specification of RBAC policies AC D, C (Kuhlmann et al., 2013)
SI* Asset & Trust 2013 identify insider threats TM R (Paci et al., 2013)
Adão et al. 2013 secure key management CR C (Adão et al., 2013)
XFPM-RBAC 2013 location-aware RABAC AC D (Unal and Çaglayan, 2013)
ZQL 2013 privacy-preserving data processing CR C (Fournet et al., 2013)
@@
BreakGlassPolicies 2014 break-glass policies in role-based access control AC R, D (Schefer-Wenzl and Strembeck, 2014a)
DelegationRBAC 2014 definition of rights delegations AC R, D (Schefer-Wenzl and Strembeck, 2014b)
DPSL 2014 facilitate specification of SOA security policies R D, RT (Wu-Lee and Hwang, 2014)
@@
Haka 2014 describe protocols and apply security policies on network traffic ID C, RT (Arkoon Network Security and ParisTech, 2014) (Fariello et al., 2016)
Jif 2014 practical information flow control IFC C (Pullicino, 2014) (Myers et al., 2022)
Sapper 2014 timing-sensitive information flow control in hardware IFC C (Li et al., 2014)
SecDSVL 2014 specification of security requirements and threat modeling R, TM R, D (Almorsy and Grundy, 2014) (Almorsy, 2013)
Wysteria 2014 secure multi party computation CR C, RT (Rastogi et al., 2014) (Rastogi et al., 2017)
XML Security DSL 2014 simplifying the use of XML security mechanisms CR C (Schefer-Wenzl and Feiertag, 2014)
AIDD 2015 rule-based attack description and response language ID, TM C (Souissi, 2015)
KEYRASE 2015 secure information erasure in software IFC, C C (Askarov et al., 2015)
SecPDL 2015 distributed security policy self-management AC D (Ma et al., 2015)
PPLv2 2015 access and usage control AC C (Cerbo et al., 2015; Trabelsi et al., 2011)
@@
SecVerilog 2015 timing-sensitive information flow control in hardware IFC C (Zhang et al., 2015) (Ferraiuolo, 2015)
STS 2015 modeling and reasoning about security requirements R R, D (Paja et al., 2015) (Giorgini et al., 2020)
SysML-Sec 2015 security requirements and identification of possible attacks R, TM R, D, C (Roudier and Apvrille, 2015)
GemRBAC-DSL 2016 expressive specification of RBAC AC D, RT (Fadhel et al., 2016) (Fadhel, 2017)
Lifty 2016 automated repair for information flow control IFC C, RT (Polikarpova et al., 2020)
@@
LJGS 2016 static information flow control supported by dynamic checks IFC C (Fennell and Thiemann, 2016a) (Fennell and Thiemann, 2016b)
pwnPr3d 2016 attack graph generation TM D (Johnson et al., 2016) (Pwn, 2017)
IoTsec 2017 specification of security requirements R R, D (Robles-Ramirez et al., 2017)
@@
Jasmin 2017 efficient, safe, secure and correct cryptographic implementation CR C (Almeida et al., 2017) (Laporte et al., 2024)
Low* 2017 efficient, safe, secure and correct cryptographic implementation CR C (Protzenko et al., 2017) (Martínez et al., 2024)
ProScript 2017 secure cryptographic protocol code CR C (Kobeissi et al., 2017) (Kobeissi et al., 2022)
Vale 2017 secure and efficient cryptographic implementation CR C (Bond et al., 2017) (Baumann et al., 2024)
SRESOFL 2017 formal specification of security requirements R R, D, C (Emeka and Liu, 2017)
STREAMS 2017 specify information flow control policies for hardware IFC C (Kwon et al., 2017)
ChiselFlow 2018 timing-sensitive information flow control in hardware IFC C (Ferraiuolo et al., 2018)
IoT Language 2018 threat modeling and risk assessment TM D (Beaulaton et al., 2018) (Beaulaton, 2018)
LOCKS 2018 formal and convenient security goal specification R, TM R, D, C (Kumar et al., 2018)
LUCON 2018 data-centric security policies for distributed systems IFC C (Schütte and Brost, 2018)
MAL 2018 threat modeling and attack simulation TM D, T (Katsikeas et al., 2019; Johnson et al., 2018; Widel et al., 2023) (Wällstedt et al., 2024)
FaCT 2019 timing side-channel secure cryptography CR C (Cauligi et al., 2017, 2019) (Soeller et al., 2019)
Flamio 2019 coarse-grained, dynamic, information-flow control IFC C (Pedersen and Chong, 2019) (fla, 2018)
IDS-DL 2019 visual description of intrusion detection systems ID D (Amine et al., 2019)
UML-SR 2019 specification of low-level security requirements R R, D (Mohsin and Khan, 2019)
Cao et al. 2019 access control between system components AC D (Cao et al., 2019)
Zhao et al. 2019 integration of functional and security models R D (Zhao et al., 2019)
QIF-RTL 2019 quantitative information flow control in hardware IFC C (Guo et al., 2019; Reimann et al., 2021) (Guo and Jin, 2019)
Zee 2019 information flow control for software systems with RTE IFC C (Vorreiter Pedersen and Askarov, 2019) (zee, 2019)
DynaMo 2020 policy analysis and automated source code generation AC C (Amthor and Schlegel, 2020)
Legal-GRL 2020 legal compliance requirements R, AC R, D, C, T (Sartoli et al., 2020)
MACasync 2020 secure information flow with respect to concurrency IFC C (Tom’e Corti nas et al., 2020)
REAP 2020 specify structure of security policies AC D (Amthor and Schlegel, 2020)
@@
Co-Inflow 2021 dynamic, coarse-grained information flow control IFC C (Xiang and Chong, 2021) (Xiang, 2021)
JE 2021 language support for TEE integration IFC C (Oak et al., 2021) (Oak et al., 2022)
SSEL 2021 specification of secure SoC component communication R R, D, C, T (Raj et al., 2021)
Hamraz 2022 information flow control in distributed system IFC C (Li et al., 2022)
IFL (IFCIL) 2022 information flow control for SELinux IFC C (Ceragioli et al., 2022)
Before examining the identified security DSLs in detail, to get an understanding of the background of the examined security DSLs, we provide an overview of the DSLs, the scientific communities from which they originate, and how many of them are still actively maintained and accessible.
Summary of the identified DSLs
Tables 1 and 2 show the security DSLs identified in the systematic literature review as well as the corresponding references. While DSLs, such as SELinks/FABLE, SecureUML, or CORAS, were presented in multiple papers that we found in our literature review, two papers contain two DSLs in a single paper (DynaMo & REAP (Amthor and Schlegel, 2020) and XACML & SAML (Ferdous and Poet, 2014)). The abbreviation of the Security Policy Description Language (Ma et al., 2015) conflicts with the abbreviation of the Simplified Policy Description Language (Nakamura et al., 2009), we therefore refer to the former with SecPDL and only the latter with SPDL. As visualized in Figure 3, the DSLs identified in our literature review were published between 1997 and 2022. It is clearly visible that the number of published DSLs increases relatively steadily until a peak in the publication of security DSLs is reached in 2013. Thereafter, the number of newly published security DSLs decreases steadily. This raises the hypothesis that security DSLs have been published for most relevant security aspects. We will explore this hypothesis in the detailed analysis of the security DSLs.
Background of the DSLs

### v3

@@ 3.2. DSL Analysis
Our in-depth analysis of the identified security DSLs followed a four-step procedure.
1. Definition of aspects: For each research question, we defined relevant aspects according to which we analyzed the identified security DSLs. We outlined these aspects already when discussing the research questions in Sec. 1 and provide detailed reasoning for each of them in Sec. 4.
2. Free-text responses: To describe each DSL according to these aspects, we first searched the corresponding papers from our dataset for relevant information. If we did not find the information in the paper, or if the aspect was open-ended (e.g., whether there are practical applications of the DSL), we searched for further publications on the DSL and for websites or documentation of the DSL. In doing so, we did not limit ourselves to academic literature for analyzing the security DSLs. In this first step, we collected short free-text responses to each aspect. To this end, we divided the selected papers between the first two authors of this paper. While collecting the information on the security DSLs, the two authors were continuously synchronizing to discuss interpretations of single aspects for specific DSLs and to avoid divergence in the collected information.
3. Categorization: After collecting information for each DSL, we discussed all captured free texts and derived categories from them. Since not all captured information was easily mappable to the identified categories, the categorization was followed by a data sanitation step in which the first two authors this time divided the individual aspects between them and went through all captured free text answers as well as the full papers again to assign every DSL, depending on the aspect, one or multiple of the identified categories. Again, we supported this procedure with continuous discussions between the first and second author.
4. Statistical analysis: Finally, after classifying all DSLs according to the categories identified for each aspect, we created statistics on the distribution of categories per aspect, including correlations among them. While discussed below, our supplementary material (Krausz et al., 2025) provides further details.
@@ 4. Systematic Literature Review
Our SLR identified 120 security DSLs. We now discuss the results of our in-depth analysis per RQ.
@@
ASL 1997 generic modeling of access control policies AC D, C (218)
Bro (Zeek) 1999 detection of network attackers by passive monitoring ID C, RT (260) (Sharma et al., 2024)
AdeLe 2001 model a database of known attack scenarios TM, ID C, RT (264)
Ponder 2001 generic modeling of access control policies AC D, C (226)
SPL 2001 single description of global security policies AC D, C (219)
@@
CORAS 2002 risk assessment based on semiformal system model TM R, D (300; 301) (SINTEF, 2014)
@@
SecureUML 2002 modeling access control policies and their integration AC D, C, RT (76; 202; 267; 289) (Doser, 2007)
STATL 2002 intrusion detection of known attack scenarios ID C, RT (209)
UMLsec 2002 planning of secure system design R D (187; 347) (Jürjens et al., 2024)
FlowCAML 2003 secure information flow language IFC C (238) (Simonet, 2003)
uniframe 2003 security contracts for component descriptions AC D (291)
Ray et al. 2003 composition of access control frameworks AC D (270)
Rei 2003 policies applicable to pervasive environments AC C, RT (227; 228)
XACML 2003 common language for expressing security policies AC C, RT (126; 229; 230) (OASIS, 2021)
ESRML 2004 enterprise information security requirements for IS0-17799 R, AC R (220)
SAML 2005 exchanging authentication and authorization data AC C, RT (126) (355)
X-GTRBAC 2005 expressive specification of RBAC AC D (279; 280; 281; 285)
WS-CoL 2005 policy assertions of user requirements for web services R C, RT (275)
UMLintr 2006 specification of intrusion detection ID R, D, C, RT (205)
WS-SecurityPolicy 2006 security constraints and requirements for web services R C, RT (272) (274)
AsmLSec 2007 specifying attack scenarios ID T (208)
OSL 2007 specify usage control policies AC D, C (324)
SECTET 2007 policies for secure interorganizational workflows R, AC R, D (244; 245; 246; 247; 290)
SecureTROPOS 2007 identification and planning of security requirements R R, D (210; 211; 212) (Mouratidis, 2015)
SMCL 2007 secure multi party computation CR C (352)
@@
Rodríguez et al. 2007 embed security requirements into business processes R R, D (257)
WS-Policy4MASC 2007 specification of monitoring and control policies R D, C (273)
CA-RBAC 2008 context-based access control AC C, RT (286)
SELinks/FABLE 2008 enforcement of security policies in web applications AC, IFC C (83; 200) (Corcoran et al., 2008)
SFDL 2.0 2008 secure multi party computation CR C, RT (349; 348) (Tang and others, 2015)
Breu et al. 2008 guide security investment decisions TM R, D (333)
USE 2008 specify and verify security policies AC D, C, RT (292)
Cryptol 2009 functional correctness of cryptography CR C (309) (Galois Inc., 2024)
Fabric 2009 secure information flow in distributed systems IFC C (239; 240) (Liu et al., 2019)
Nego-UCONABC 2009 access control for cloud services AC D (306)
ReAlSec 2009 identify countermeasures for known attack patterns TM T, DP (214)
S-Promela 2009 generic security policies AC C (215)
Security Policy Model 2009 generating security policies from security requirements R, AC R, D, C, RT (256; 259; 293)
SPDL 2009 simplifying the specification of SELinux security policies AC C, RT (243) (Ynakam, 2008)
Pavlich-Mariscal et al. 2010 access control policies AC D, C (269)
ADTrees 2010 specify and analyze attack-defense scenarios TM D (340)
ASLan++ 2010 security goals of distributed system designs AC D (338)
CapDL 2010 capability-based access control AC D, RT (77) (McLeod et al., 2024)
CRePE 2010 fine-grained security policies for android applications AC C, RT (250)
SecPAL 2010 simple but expressive, declarative authorization language AC C (316)
SSMF 2010 identification of possible attack vectors TM R (235)
TASTYL 2010 secure multi party computation CR C, RT (353)
X-Policy 2010 dynamic execution permissions for collaborative systems AC D (315)
Caisson 2011 secure information flow language IFC C (236)
cPLC 2011 generate secure code for cryptographic two-party protocols CR D, C (90)
Asnar et al. 2011 identification of permission-related threats R, TM R (233)
UACML 2011 access control modeling language independent of model AC D (71)
Bain et al. 2011 high-level code for secure execution platforms (MPC & FHE) CR C (196)
VALID 2011 security goals for virtualized environments R DP, RT (224)
PTaCL 2012 security policies based on attribute-based access control AC D (346)
Hoisl et al. 2012 secure object flow in business process IFC D (192)
STRoBAC 2012 spatial temporal role-based access control AC D, C, RT (283)
@@
TTCN-3 2012 generation of penetration tests with attack patterns TM T (213) (ETSI, 2023)
Mitchell et al. 2012 secure execution platform based on MPC or FHE CR C (102)
BCDSDL 2013 network availability with respect to botnet attacks IDE D (68)
context-aware RBAC 2013 context-aware access control policies AC D (193)
CIL 2013 specification of access control policies for Android AC C, RT (25) (SELinux Project, 2023)
@@
CySeMoL 2013 probabilistic vulnerability assessment TM D, C (133; 136) (KTH Network and Systems Engineering Group, 2020)
DJS 2013 solution for untrusted code in javascript CR C (46)
FAL 2013 secure logging configurations CR C, RT (222)
Paragon 2013 practical information flow control IFC C (242) (Broberg et al., 2013)
Table 2. Security DSLs identified in the SLR (Part 2/2)
@@
RBAC DSL 2013 specification of RBAC policies AC D, C (104)
SI* Asset & Trust 2013 identify insider threats TM R (105)
Adão et al. 2013 secure key management CR C (109)
XFPM-RBAC 2013 location-aware RABAC AC D (166)
ZQL 2013 privacy-preserving data processing CR C (96)
@@
BreakGlassPolicies 2014 break-glass policies in role-based access control AC R, D (194)
DelegationRBAC 2014 definition of rights delegations AC R, D (195)
DPSL 2014 facilitate specification of SOA security policies R D, RT (169)
@@
Haka 2014 describe protocols and apply security policies on network traffic ID C, RT (266) (Fariello et al., 2016)
Jif 2014 practical information flow control IFC C (188) (Myers et al., 2022)
Sapper 2014 timing-sensitive information flow control in hardware IFC C (134)
SecDSVL 2014 specification of security requirements and threat modeling R, TM R, D (11) (Almorsy, 2013)
Wysteria 2014 secure multi party computation CR C, RT (37) (Rastogi et al., 2017)
XML Security DSL 2014 simplifying the use of XML security mechanisms CR C (175)
AIDD 2015 rule-based attack description and response language ID, TM C (69)
KEYRASE 2015 secure information erasure in software IFC, C C (63)
SecPDL 2015 distributed security policy self-management AC D (158)
PPLv2 2015 access and usage control AC C (118; 323)
@@
SecVerilog 2015 timing-sensitive information flow control in hardware IFC C (132) (Ferraiuolo, 2015)
STS 2015 modeling and reasoning about security requirements R R, D (248) (Giorgini and others, 2020)
SysML-Sec 2015 security requirements and identification of possible attacks R, TM R, D, C (191)
GemRBAC-DSL 2016 expressive specification of RBAC AC D, RT (225) (Fadhel, 2017)
Lifty 2016 automated repair for information flow control IFC C, RT (241)
@@
LJGS 2016 static information flow control supported by dynamic checks IFC C (139) (Fennell and Thiemann, 2016)
pwnPr3d 2016 attack graph generation TM D (91) (92)
IoTsec 2017 specification of security requirements R R, D (206)
@@
Jasmin 2017 efficient, safe, secure and correct cryptographic implementation CR C (198) (Laporte and others, 2024)
Low* 2017 efficient, safe, secure and correct cryptographic implementation CR C (199) (Martínez et al., 2024)
ProScript 2017 secure cryptographic protocol code CR C (81) (Kobeissi et al., 2022)
Vale 2017 secure and efficient cryptographic implementation CR C (82) (Baumann et al., 2024)
SRESOFL 2017 formal specification of security requirements R R, D, C (5)
STREAMS 2017 specify information flow control policies for hardware IFC C (43)
ChiselFlow 2018 timing-sensitive information flow control in hardware IFC C (299)
IoT Language 2018 threat modeling and risk assessment TM D (146) (Beaulaton, 2018)
LOCKS 2018 formal and convenient security goal specification R, TM R, D, C (145)
LUCON 2018 data-centric security policies for distributed systems IFC C (101)
MAL 2018 threat modeling and attack simulation TM D, T (17; 22; 189) (Wällstedt and others, 2024)
FaCT 2019 timing side-channel secure cryptography CR C (29; 197) (Soeller et al., 2019)
Flamio 2019 coarse-grained, dynamic, information-flow control IFC C (88) (34)
IDS-DL 2019 visual description of intrusion detection systems ID D (28)
UML-SR 2019 specification of low-level security requirements R R, D (2)
Cao et al. 2019 access control between system components AC D (18)
Zhao et al. 2019 integration of functional and security models R D (182)
QIF-RTL 2019 quantitative information flow control in hardware IFC C (12; 190) (Guo and Jin, 2019)
Zee 2019 information flow control for software systems with RTE IFC C (21) (122)
DynaMo 2020 policy analysis and automated source code generation AC C (157)
Legal-GRL 2020 legal compliance requirements R, AC R, D, C, T (80)
MACasync 2020 secure information flow with respect to concurrency IFC C (249)
REAP 2020 specify structure of security policies AC D (157)
@@
Co-Inflow 2021 dynamic, coarse-grained information flow control IFC C (54) (Xiang, 2021)
JE 2021 language support for TEE integration IFC C (6) (Oak and others, 2022)
SSEL 2021 specification of secure SoC component communication R R, D, C, T (1)
Hamraz 2022 information flow control in distributed system IFC C (53)
IFL (IFCIL) 2022 information flow control for SELinux IFC C (13)
We first provide an overview of the security DSLs, their background, the scientific communities from which they originate, and how many of them are still actively maintained and accessible. Tables 1 and 2 show the identified DSLs with their references. Two DSLs share the same abbreviation, so we refer to the former (158) as SecPDL and the latter (243) as SPDL. Figure 3 shows that the DSLs were published between 1997 and 2022, and that the number of published DSLs increases steadily until a peak is reached in 2013. We believe that DSLs have been published for most relevant security aspects—a hypothesis we explored in our detailed analysis discussed below.
4.1.1. Background of the DSLs